Due to erratic sales of its sole product—a high-capacity battery for laptop computers—PEM, Inc., has been experiencing financial difficulty for some time. The company’s contribution format income statement for the most recent month is given below:

  

Sales (12,800 units × $20 per unit) $ 256,000
Variable expenses 128,000
Contribution margin 128,000
Fixed expenses 143,000
Net operating loss $ (15,000 )

Required:

1. Compute the company’s CM ratio and its break-even point in unit sales and dollar sales.

2. The president believes that a $6,400 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in an $85,000 increase in monthly sales. If the president is right, what will be the increase (decrease) in the company’s monthly net operating income?

3. Refer to the original data. The sales manager is convinced that a 10% reduction in the selling price, combined with an increase of $35,000 in the monthly advertising budget, will double unit sales. If the sales manager is right, what will be the revised net operating income (loss)?

4. Refer to the original data. The Marketing Department thinks that a fancy new package for the laptop computer battery would grow sales. The new package would increase packaging costs by $0.50 per unit. Assuming no other changes, how many units would have to be sold each month to attain a target profit of $4,600?

5. Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $52,000 each month.

a. Compute the new CM ratio and the new break-even point in unit sales and dollar sales.

b. Assume that the company expects to sell 20,500 units next month. Prepare two contribution format income statements, one assuming that operations are not automated and one assuming that they are. (Show data on a per unit and percentage basis, as well as in total, for each alternative.)

c. Would you recommend that the company automate its operations (Assuming that the company expects to sell 20,500)?
Answer #1

1.CM Ratio = Contribution Margin/Sales

= 128000/256,000

= 50%

Break even point in unit sales = Fixed costs/Contribution Margin per unit

= 143000/10

= 14,300 units

Dollar sales = 143000/50%

= $286,000

2.Increase in Net Operating Income = Increase in Contribution Margin – Increase in cost

= 85,000*50% - 6400

= $36,100

3.

Contribution format Income Statement

Sales 25600*18

460,800

Less: Variable expenses

256000

Contribution Margin

204,800

Less: Fixed Expenses

178000

Net Operating Income

26,800

4.Target Profit = $4600

Fixed costs = 143000

Target Contribution Margin = $147600

Units required to be sold = 147600/(20-10-0.5)

= 15,536.84 units

5.CM Ratio = (20-7)/20 = 65%

Break even point in unit sales = (143000+52,000)/13 = 15,000 units

Dollar sales = 195000/65%

= $300,000

b.

Without Automation

With Automation

Per unit

Total

%

Per unit

Total

%

Sales

20

410,000

100.00%

20

410,000

100.00%

Variable expenses

10

205,000

50.00%

7

143,500

35.00%

Contribution Margin

10

205,000

50.00%

13

266,500

65.00%

Fixed Expenses

6.97561

143,000

34.88%

9.5122

195,000

47.56%

Net operating income

3.02439

62,000

15.12%

3.4878

71,500

17.44%

Yes, should automate